# On endosplit $p$-permutation resolutions and Broué’s conjecture for $p$-solvable groups

### v2

@@ 3. Endosplit $p$-permutation resolutions and relatively endotrivial complexes
For this section, $V$ is an absolutely $p$-divisible $kG$-module, possibly 0. We review the notions of endosplit $p$-permutation resolutions, first introduced by Rickard in [Ric96], a class of “invertible” endotrivial $p$-permutation resolutions, endotrivial complexes, introduced in [Mil24], and their relative counterpart, introduced in [Mil25c].
@@
[Ric96, Section 7] Let $M$ be a $kG$-module. An endosplit $p$-permutation resolution of $M$ is a bounded chain complex $C_{M}$ of $p$-permutation $kG$-modules such that the homology of $C_{M}$ is concentrated in a single degree $i\in\mathbb{Z}$, $H_{i}(C_{M})\cong M$, and the complex $\operatorname{End}_{k}(C_{M})\cong C_{M}^{*}\otimes_{k}C_{M}$ is split as a chain complex of $kG$-modules.
@@
It is straightforward to see via the Künneth formula that a bounded chain complex $C$ of $p$-permutation modules is an endosplit $p$-permutation resolution (for some $kG$-module) if and only if $C^{*}\otimes_{k}C\simeq N[0]$ for some $kG$-module $N$, since split complexes are homotopy equivalent to their homology regarded as a chain complex with zero differential, and the Künneth formula implies $N\cong H_{i}(C)\otimes_{k}H_{i}(C)^{*}$ for a unique integer $i\in\mathbb{Z}$. Observe that since $N[0]$ a direct summand of $C^{*}\otimes_{k}C$, a complex of $p$-permutation modules, $N$ is $p$-permutation, so $H_{i}(C)$ is endo-$p$-permutation. In other words, if a $kG$-module has an endosplit $p$-permutation resolution, it is endo-$p$-permutation.
@@
Let $C$ be a bounded chain complex of $p$-permutation $kG$-modules. The following are equivalent:
@@
Now, suppose condition (b) holds. Suppose for contradiction that $C^{*}\otimes_{k}C\not\simeq M[0]$, then we may write $C^{*}\otimes_{k}C\simeq D$ where $D$ is assumed to have no contractible summands. Since $C$ is bounded, there exists a unique positive natural number $i>0$ for which $D_{i}\neq 0$ and $D_{j}=0$ for all $j>i$. Since $C(P)$ is either contractible or has homology contained in exactly one degree, $C(P)^{*}\otimes_{k}C(P)$ has nonzero homology concentrated in degree 0, possibly also zero. The same holds for $D(P)$, as it is a direct summand of $(C^{*}\otimes_{k}C)(P)\cong C(P)^{*}\otimes_{k}C(P).$ In particular, $d_{i}(P):D_{i}(P)\to D_{i-1}(P)$ is injective for all $P\in s_{p}(G)$, so $d_{i}:D_{i}\to D_{i-1}$ is split injective by [Lin18a, Proposition 5.8.11]. Therefore $D$ has a contractible direct summand, a contradiction. Thus, $C^{*}\otimes_{k}C\simeq M[0]$ for some $kG$-module $M$. The Künneth formula implies there exists a unique $i\in\mathbb{Z}$ for which $H_{i}(C)\neq 0$. We conclude $C$ is an endosplit $p$-permutation resolution. ∎
@@
[Lin18b, Theorem 7.11.2] Let $M$ be a $kG$-module with endosplit $p$-permutation resolution resolution $C_{M}$. For any subgroup $H\leq G$, the complex $\operatorname{End}_{kH}(C)$ has homology concentrated in degree zero, isomorphic to $\operatorname{End}_{kH}(M)$, and there is a $k$-algebra isomorphism
$\rho:\operatorname{End}_{K^{b}({}_{kH}\mathbf{mod})}(C)\cong\operatorname{End}_{kH}(M)$
@@
[Ric96, Proposition 7.1] Let $G$ be a finite group, and let $M$ be a $kG$-module that has an endosplit $p$-permutation resolution $X_{M}$. Then $M$ lifts to a unique $\mathcal{O}G$-module $\tilde{M}$, and $X_{M}$ lifts to a unique endosplit $p$-permutation resolution $\tilde{X}_{\tilde{M}}$ of $M$, i.e. $X_{M}\cong k\otimes_{\mathcal{O}}\tilde{X}_{\tilde{M}}$.
@@
[Mil25c, Definition 6.1] Let $C$ be a bounded chain complex of $p$-permutation $kG$-modules. $C$ is a $V$-endosplit-trivial complex if and only if $C^{*}\otimes_{k}C\simeq(k\oplus N)[0]$ for some $V$-projective $kG$-module $N$. Equivalently, $C$ is an endosplit $p$-permutation resolution of a $V$-endotrivial $kG$-module. For shorthand, we refer to $V$-endosplit-trivial complexes as simply $V$-endotrivial complexes. $V$-endotrivial complexes over $\mathcal{O}$ are defined analogously.
If $V=0$, then we have $C^{*}\otimes_{k}C\simeq k[0]$, recovering the definition of an endotrivial complex, see [Mil24, Definition 3.1]. These are the invertible objects of the tensor-triangulated category $K^{b}({}_{kG}\mathbf{triv})$, and are completely classified in [Mil25a].
@@
If $C^{*}\otimes_{k}C\simeq(k\oplus N)[0]$, $N$ is necessarily $p$-permutation. Therefore, when considering the class of $V$-endotrivial complexes, $V$ can be replaced with a suitable $p$-permutation $kG$-module $W$.
@@
[Mil25c, Propositions 6.6, 6.10] Let $C$ be a $V$-endotrivial complex.
@@
Neither of $(C_{0}^{*}\otimes_{k}C_{1}),(C_{1}^{*}\otimes_{k}C_{0})$ can contain $k[0]$ as a direct summand since they are dual, therefore one of the other two summands contains $k[0]$ as a direct summand. Without loss of generality, suppose that it is $C_{0}^{*}\otimes_{k}C_{0}$. Then $C_{0}$ is a $V$-endotrivial chain complex and the direct sum of the three other terms in the direct sum is homotopy equivalent to $N^{\prime}[0]$ for some $V$-projective $kG$-module $N^{\prime}$. Therefore, all three other terms in the direct sum are $V$-projective as chain complexes.
@@
For (b), Let $S\in\operatorname{Syl}_{p}(G)$. Since $(C^{*}\otimes_{k}C)(S)\cong k\neq 0$, $C$ contains at least one degree with Sylow vertices. Therefore, $C_{0}$ itself has Sylow vertices. ∎
@@
We impose an equivalence condition $\sim_{V}$ on the class of $V$-endotrivial chain complexes of $kG$-modules (analogous to the module-theoretic case) as follows: if $C$ and $D$ are $V$-endotrivial chain complexes, write $C\sim_{V}D$ if and only if $C$ and $D$ have isomorphic indecomposable $V$-endotrivial summands. Let $\mathcal{E}_{k}^{V}(G)$ denote the resulting set of equivalence classes. $\mathcal{E}_{k}^{V}(G)$ is an abelian group with addition defined by $[C]+[D]:=[C\otimes_{k}D]$. We have $-[C]=[C^{*}]$ and each equivalence class has a unique indecomposable representative.
@@
Let $C$ be a bounded chain complex of $p$-permutation $kG$-modules. The following are equivalent.
@@
    $C$ is an endosplit $p$-permutation resolution of a $V$-endotrivial $kG$-module.
@@
First, suppose $C$ is a $V$-endotrivial complex. Since $C^{*}\otimes_{k}C\simeq(k\oplus N)[0]$ for some $V$-projective $kG$-module $N$,
@@
It follows by the Künneth formula that $C(P)$ also has homology concentrated in a single degree, $h_{C}(P)$, and if $V(P)=0$, then since $N$ is $V$-projective, $N=0$. This shows (c).
@@
$\dim_{k}H_{h_{C}(P)}(C(P))=1$ for all $P\not\in\mathcal{X}_{V}$, $C(P)^{*}\otimes_{k}C(P)\cong(C^{*}\otimes_{k}C)(P)\simeq k[0]$. On the other hand, since $C$ is an endosplit $p$-permutation resolution, $C^{*}\otimes_{k}C\simeq(M^{*}\otimes_{k}M)[0]$. Thus $M^{*}\otimes_{k}M\cong L$, where $L$ is a $p$-permutation $kG$-module for which $L(P)\cong k$ for any $P\not\in\mathcal{X}_{V}$.
Since $V$ is $p$-permutation, $V$-projectivity is equivalent to $\mathcal{X}_{V}$-projectivity, where $\mathcal{X}_{V}$ is the set of $p$-subgroups for which $V(P)\neq 0$. A $p$-permutation module $N$ is $\mathcal{X}_{V}$-projective if and only if $N(P)=0$ for all $P\not\in\mathcal{X}_{V}$, since $\mathcal{X}_{V}$ is downward closed as a poset. Therefore, $L=L_{0}\oplus L_{1}$, where $L_{0}$ is a $p$-permutation $kG$-module with vertex $S$ and $L_{1}$ is $V$-projective. Since $L_{0}$ has vertex $S$ and $L_{0}(S)\cong k$, $L_{0}\cong k$ since $L_{0}(S)$ is the Green correspondent of $L_{0}$ [Lin18a, Proposition 5.10.5]. We conclude that $M$ is $V$-endotrivial, as desired. ∎
@@
[Mil25c, Definition 9.6] If $C$ is a $V$-endotrivial complex, the values $h_{C}(P)$ in the previous proposition define an integer-valued class function on the $p$-subgroups of $G$, $h_{C}:s_{p}(G)\to\mathbb{Z}$. We denote the set of such class functions by $CF(G,p)$, call the functions $h_{C}$ the h-marks of $C$, and say $h_{C}(P)$ is the h-mark of $C$ at $P$.
It is straightforward to verify that the map $h:\mathcal{E}_{k}^{V}(G)\to CF(G,p)$ is a well-defined group homomorphism. We call $h$ the h-mark homomorphism, as the values $h_{C}(P)$ are analogous to the marks of a $G$-set. If $X$ is a $G$-set, then $X$ is uniquely determined by the marks $|X^{H}|$ where $H$ runs over all conjugacy class representatives of subgroups of $G$.
@@
Let $T_{k,V}(G,S)\leq T_{k,V}(G)$ denote the subgroup of $T_{k,V}(G)$ consisting of $p$-permutation $V$-endotrivial $kG$-modules. Equivalently, if $S\in\operatorname{Syl}_{p}(G)$, $T_{k,V}(G,S)$ is the kernel of the restriction map $\operatorname{Res}^{G}_{S}:T_{k,V}(G)\to T_{k,V}(S)$. $T_{k,V}(G,S)$ is a finite group.
@@
[Mil25c, Theorem 9.8], [Mil25a, Theorem 2.14] $h:\mathcal{E}_{k}^{V}(G)\to CF(G,p)$ has kernel the torsion subgroup of $\mathcal{E}_{k}^{V}(G)$, which is identified with
$T_{k,V}(G,S)$ via the inclusion $[M]\mapsto[M[0]]$. In particular, if $G$ is a $p$-group, $h$ is injective. If $V=V(\mathcal{F}_{G})$, $h$ is surjective.
@@
The kernel of $h$ consists of (equivalence classes of) $V$-endotrivial chain complexes $C$ for which $C(P)$ has nonzero homology only at degree 0 for any $p$-subgroup $P$. Suppose for contradiction that there exists an indecomposable $V$-endotrivial chain complex $C$ with $[C]\in\ker h$ such that $C\not\cong M[0]$ for some $kG$-module $M$. Without loss of generality suppose that there exists $i>0$ for which $C_{i}\neq 0$, and $C_{j}=0$ for all $j>i$. Then for all $p$-subgroups $P$, $d_{i}(P):C_{i}(P)\to C_{i-1}(P)$ is injective, hence $d_{i}$ is split injective, a contradiction to $C$ being indecomposable. Thus if $C$ is indecomposable and $[C]\in\ker h$, $C\cong M[0]$ for a necessarily $p$-permutation $V$-endotrivial $kG$-module $M$. Thus, $\ker h\cong T_{k,V}(G,S)$. Since $T_{k,V}(G,S)$ is finite, $\ker h$ is torsion, and any $[C]\not\in\ker h$ clearly cannot be torsion since for any $n\in\mathbb{N}$, $n\cdot h_{C}$ outputs a nonzero value.
@@
under $h$ forms a diagonal basis of $CF(G,p)$, hence $h$ is surjective, and in fact, the above set is a basis for $\mathcal{E}_{k}^{V(\mathcal{F}_{G})}(G)$. ∎
@@
It is clear that every $kG$-module that has an endosplit $p$-permutation resolution is endo-$p$-permutation, but, in general, not every endo-$p$-permutation module has an endosplit $p$-permutation resolution. Mazza proved in [Maz03] when $p$ is odd and $G$ is a $p$-group, every endopermutation $kG$-module (hence endo-$p$-permutation $kG$-module) has an endosplit $p$-permutation resolution. On the other hand, if $G$ is a $2$-group, not every endopermutation $kG$-module has an endosplit $p$-permutation resolution; for instance, the unique endotrivial $kQ_{8}$-module that does not descend to $\mathbb{F}_{2}$ has no endosplit $p$-permutation resolution.
@@
    $p$-permutation resolution $C$ as follows. Let $\mathcal{X}_{C}$ denote the set of all $p$-subgroups of $G$ for which $C(P)\not\simeq 0$. It is easy to see that $\mathcal{X}_{C}$ forms a downward closed sub-poset of $s_{p}(G)$ which is closed under $G$-conjugation. For each $P\in\mathcal{X}_{C}$, write $h_{C}(P)$ for the unique integer for which $H_{h_{C}(P)}(C(P))\neq 0$. This defines a class function on $\mathcal{X}_{V}$, $h_{C}:\mathcal{X}_{C}\to CF(G,\mathcal{X}_{C})$.
@@
    We have that $C_{M}$ is a $V(\mathcal{F}_{G})$-endotrivial chain complex; we say a complex isomorphic to a shift of a complex in the above form is a chain tensor of relative syzygies. Given any chain tensor of relative syzygies $C$, there exists a unique corresponding module tensor of relative syzygies, obtained by taking the nonzero homology $H_{h_{C}(1)}(C)$. Conversely, it is clear from the definition that given any module $M$ tensor of relative syzygies, there exists some chain tensor of relative syzygies $C_{M}$ resolving $M$, however $C_{M}$ need not be unique.
@@
We are about to characterize all $kG$-modules which have endosplit $p$-permutation resolutions, as well as all endosplit $p$-permutation resolutions themselves. This may be seen as a classification theorem, as there is no group structure that parameterizes all endo-$p$-permutation $kG$-modules (which have endosplit $p$-permutation resolutions) or all endosplit $p$-permutation resolutions, due to a lack of invertibility of some objects, in a sense.
Similar to how capped endopermutation modules can be thought of as the invertible endopermutation modules, and thus can be given a group structure, the $V(\mathcal{F}_{G})$-endotrivial complexes can be thought of as the invertible, or “capped,” endosplit $p$-permutation resolutions. The following theorem morally states that every endo-$p$-permutation module (resp. endosplit $p$-permutation resolution) can be built from invertible endo-$p$-permutation modules (resp. endosplit-$p$-permutation resolutions), similar to how all endopermutation modules for $p$-groups arise from capped endopermutation modules (see [Dad78, Theorem 6.6]).

### v3

@@ 3. Endosplit $p$-permutation resolutions and relatively endotrivial complexes
For this section, $V$ is an absolutely $p$-divisible kGkG-module, possibly 0. We review the notions of endosplit $p$-permutation resolutions, first introduced by Rickard in [R96], a class of “invertible” endotrivial $p$-permutation resolutions, endotrivial complexes, introduced in [M24a], and their relative counterpart, introduced in [M24b].
@@
[R96, Section 7] Let $M$ be a $kG$-module. An endosplit $p$-permutation resolution of $M$ is a bounded chain complex $C_{M}$ of $p$-permutation kGkG-modules such that the homology of $C_{M}$ is concentrated in a single degree $i\in\mathbb{Z}$, $H_{i}(C_{M})\cong M$, and the complex $\operatorname{End}_{k}(C_{M})\cong C_{M}^{*}\otimes_{k}C_{M}$ is split as a chain complex of $kG$-modules.
@@
It is straightforward to see via the Künneth formula that a bounded chain complex $C$ of $p$-permutation modules is an endosplit $p$-permutation resolution (for some $kG$-module) if and only if $C^{*}\otimes_{k}C\simeq N[0]$ for some $kG$-module NN, since split complexes are homotopy equivalent to their homology regarded as a chain complex with zero differential, and the Künneth formula implies $N\cong H_{i}(C)\otimes_{k}H_{i}(C)^{*}$ for a unique integer $i\in\mathbb{Z}$. Observe that since $N[0]$ a direct summand of $C^{*}\otimes_{k}C$, a complex of $p$-permutation modules, $N$ is $p$-permutation, so $H_{i}(C)$ is endo-$p$-permutation. In other words, if a $kG$-module has an endosplit $p$-permutation resolution, it is endo-$p$-permutation.
@@
Let $C$ be a bounded chain complex of $p$-permutation kGkG-modules. The following are equivalent:
@@
Now, suppose condition (b) holds. Suppose for contradiction that $C^{*}\otimes_{k}C\not\simeq M[0]$, then we may write $C^{*}\otimes_{k}C\simeq D$ where $D$ is assumed to have no contractible summands. Since $C$ is bounded, there exists a unique positive natural number $i>0$ for which $D_{i}\neq 0$ and $D_{j}=0$ for all $j>i$. Since $C(P)$ is either contractible or has homology contained in exactly one degree, $C(P)^{*}\otimes_{k}C(P)$ has nonzero homology concentrated in degree 0, possibly also zero. The same holds for $D(P)$, as it is a direct summand of $(C^{*}\otimes_{k}C)(P)\cong C(P)^{*}\otimes_{k}C(P).$ In particular, $d_{i}(P)\colon D_{i}(P)\to D_{i-1}(P)$ is injective for all $P\in s_{p}(G)$, so $d_{i}\colon D_{i}\to D_{i-1}$ is split injective by [L181, Proposition 5.8.11]. Therefore $D$ has a contractible direct summand, a contradiction. Thus, $C^{*}\otimes_{k}C\simeq M[0]$ for some $kG$-module $M$. The Künneth formula implies there exists a unique $i\in\mathbb{Z}$ for which $H_{i}(C)\neq 0$. We conclude $C$ is an endosplit $p$-permutation resolution. ∎
@@
[L182, Theorem 7.11.2] Let $M$ be a $kG$-module with endosplit $p$-permutation resolution resolution $C_{M}$. For any subgroup $H\leq G$, the complex $\operatorname{End}_{kH}(C)$ has homology concentrated in degree zero, isomorphic to $\operatorname{End}_{kH}(M)$, and there is a $k$-algebra isomorphism
$\rho_{H}\colon\operatorname{End}_{K^{b}({}_{kH}\mathbf{mod})}(C)\cong\operatorname{End}_{kH}(M)$
@@
[R96, Proposition 7.1] Let $G$ be a finite group, and let $M$ be a $kG$-module that has an endosplit $p$-permutation resolution $X_{M}$. Then there exists a unique lift of $M$ to a $\mathcal{O}G$-module $\tilde{M}$ with an endosplit $p$-permutation resolution $\tilde{X}_{M}$ lifting $X_{M}$, i.e. $X_{M}\cong k\otimes_{\mathcal{O}}\tilde{X}_{M}$.
Note that $M$ may not lift uniquely to $\mathcal{O}$ - rather, the proposition guarantees the existence of a unique lift to $\mathcal{O}$ with an endosplit $p$-permutation resolution.
@@
[M24b, Definition 6.1] Let $C$ be a bounded chain complex of $p$-permutation kGkG-modules. $C$ is a $V$-endosplit-trivial complex if and only if $C^{*}\otimes_{k}C\simeq(k\oplus N)[0]$ for some $V$-projective kGkG-module NN. Equivalently, $C$ is an endosplit $p$-permutation resolution of a $V$-endotrivial kGkG-module. For shorthand, we refer to $V$-endosplit-trivial complexes as simply $V$-endotrivial complexes. $V$-endotrivial complexes over $\mathcal{O}$ are defined analogously.
If $V=0$, then we have $C^{*}\otimes_{k}C\simeq k[0]$, recovering the definition of an endotrivial complex, see [M24a, Definition 3.1]. These are the invertible objects of the tensor-triangulated category $K^{b}({}_{kG}\mathbf{triv})$, and are completely classified in [M24c].
@@
If $C^{*}\otimes_{k}C\simeq(k\oplus N)[0]$, $N$ is necessarily $p$-permutation. Therefore, when considering the class of $V$-endotrivial complexes, $V$ can be replaced with a suitable $p$-permutation kGkG-module $W$.
@@
[M24b, Propositions 6.6, 6.10] Let $C$ be a $V$-endotrivial complex.
@@
Neither of $(C_{0}^{*}\otimes_{k}C_{1}),(C_{1}^{*}\otimes_{k}C_{0})$ can contain $k[0]$ as a direct summand since they are dual, therefore one of the other two summands contains $k[0]$ as a direct summand. Without loss of generality, suppose that it is $C_{0}^{*}\otimes_{k}C_{0}$. Then $C_{0}$ is a $V$-endotrivial chain complex and the direct sum of the three other terms in the direct sum is homotopy equivalent to $N^{\prime}[0]$ for some $V$-projective kGkG-module $N^{\prime}$. Therefore, all three other terms in the direct sum are $V$-projective as chain complexes.
@@
For (b), Let $S\in\operatorname{Syl}_{p}(G)$. Since $(C^{*}\otimes_{k}C)(S)\cong k\neq 0$, $C$ contains at least one degree with Sylow vertices. From the decomposition $C\simeq C_{0}\oplus D$ shown in (a), $D(S)=0$ since $V$ is absolutely $p$-divisible. Therefore $C_{0}$ necessarily has vertex $S$, as desired. ∎
@@
We impose an equivalence condition $\sim_{V}$ on the class of $V$-endotrivial chain complexes of $kG$-modules (analogous to the module-theoretic case) as follows: if $C$ and $D$ are $V$-endotrivial chain complexes, write $C\sim_{V}D$ if and only if $C$ and $D$ have isomorphic indecomposable $V$-endotrivial summands. Let $\mathcal{E}_{k}^{V}(G)$ denote the resulting set of equivalence classes. $\mathcal{E}_{k}^{V}(G)$ is an abelian group with addition defined by $C]+[D]:=[C\otimes_{k}D]$. We have $C]=[C^{*}]$ and each equivalence class has a unique indecomposable representative.
@@
Let $C$ be a bounded chain complex of $p$-permutation kGkG-modules. The following are equivalent.
@@
    $C$ is an endosplit $p$-permutation resolution of a $V$-endotrivial kGkG-module.
@@
First, suppose $C$ is a $V$-endotrivial complex. Since $C^{*}\otimes_{k}C\simeq(k\oplus N)[0]$ for some $V$-projective kGkG-module NN,
@@
It follows by the Künneth formula that $C(P)$ also has homology concentrated in a single degree, $h_{C}(P)$, and if $V(P)=0$, then since $N$ is $V$-projective with $V$ absolutely $p$-divisible, $N(P)=0$. This shows (c).
@@
$\dim_{k}H_{h_{C}(P)}(C(P))=1$ for all $P\not\in\mathcal{X}_{V}$, $C(P)^{*}\otimes_{k}C(P)\cong(C^{*}\otimes_{k}C)(P)\simeq k[0]$. On the other hand, since $C$ is an endosplit $p$-permutation resolution, $C^{*}\otimes_{k}C\simeq(M^{*}\otimes_{k}M)[0]$. Thus $M^{*}\otimes_{k}M\cong L$, where $L$ is a $p$-permutation kGkG-module for which $L(P)\cong k$ for any $P\not\in\mathcal{X}_{V}$.
Since $V$ is $p$-permutation, $V$-projectivity is equivalent to $\mathcal{X}_{V}$-projectivity, where $\mathcal{X}_{V}$ is the set of $p$-subgroups for which $V(P)\neq 0$. A $p$-permutation module $N$ is $\mathcal{X}_{V}$-projective if and only if $N(P)=0$ for all $P\not\in\mathcal{X}_{V}$, since $\mathcal{X}_{V}$ is downward closed as a poset. Therefore, $L=L_{0}\oplus L_{1}$, where $L_{0}$ is a $p$-permutation kGkG-module with vertex $S$ and $L_{1}$ is $V$-projective. Since $L_{0}$ has vertex $S$ and $L_{0}(S)\cong k$, $L_{0}\cong k$ since $L_{0}(S)$ is the Green correspondent of $L_{0}$, see [L181, Proposition 5.10.5]. We conclude that $M$ is $V$-endotrivial, as desired. ∎
@@
[M24b, Definition 9.6] If $C$ is a $V$-endotrivial complex, the values $h_{C}(P)$ in the previous proposition define an integer-valued superclass function on the $p$-subgroups of $G$, $h_{C}\colon s_{p}(G)\to\mathbb{Z}$. We denote the set of such superclass functions by $CF(G,p)$, call the functions $h_{C}$ the $h$-marks of $C$, and say $h_{C}(P)$ is the $h$-mark of $C$ at $P$.
It is straightforward to verify that the map $h\colon\mathcal{E}_{k}^{V}(G)\to CF(G,p)$ is a well-defined group homomorphism. We call $h$ the $h$-mark homomorphism, as the values $h_{C}(P)$ are analogous to the marks of a $G$-set. That is, if $X$ is a $G$-set, then $X$ is uniquely determined by its marks $|X^{H}|$ where $H$ runs over all conjugacy class representatives of subgroups of $G$; in fact, the assignment $X\mapsto(H\mapsto|X^{H}|)$ induces an injective ring homomorphism $B(G)\to CF(G)$ (see e.g., [Bou10, Theorem 2.4.5]), where $B(G)$ denotes the Burnside ring of $G$, the split Grothendieck ring of $_{G}\mathbf{set}$.
@@
Let $T_{k,V}(G,S)\leq T_{k,V}(G)$ denote the subgroup of $T_{k,V}(G)$ consisting of $p$-permutation $V$-endotrivial kGkG-modules. Equivalently, if $S\in\operatorname{Syl}_{p}(G)$, $T_{k,V}(G,S)$ is the kernel of the restriction map $\operatorname{Res}^{G}_{S}\colon T_{k,V}(G)\to T_{k,V}(S)$. The group $T_{k,V}(G,S)$ is finite.
@@
[M24b, Theorem 9.7], [M24c, Theorem 2.12] The homomorphism $h\colon\mathcal{E}_{k}^{V}(G)\to CF(G,p)$ has kernel the torsion subgroup of $\mathcal{E}_{k}^{V}(G)$, which is identified with $T_{k,V}(G,S)$ via the inclusion $M]\mapsto[M[0]]$. In particular, if $G$ is a $p$-group, $h$ is injective. If $V=V(\mathcal{F}_{G})$, $h$ is surjective.
@@
The kernel of $h$ consists of (equivalence classes of) $V$-endotrivial chain complexes $C$ for which $C(P)$ has nonzero homology only at degree 0 for any $p$-subgroup $P$. Suppose for contradiction that there exists an indecomposable $V$-endotrivial chain complex $C$ with $C]\in\ker h$ such that $C\not\cong M[0]$ for some $kG$-module $M$. Without loss of generality suppose that there exists $i>0$ for which $C_{i}\neq 0$, and $C_{j}=0$ for all $j>i$. Then for all $p$-subgroups $P$, $d_{i}(P)\colon C_{i}(P)\to C_{i-1}(P)$ is injective, hence $d_{i}$ is split injective, a contradiction to $C$ being indecomposable. Thus if $C$ is indecomposable and $C]\in\ker h$, $C\cong M[0]$ for a necessarily $p$-permutation $V$-endotrivial kGkG-module $M$. Thus, $\ker h\cong T_{k,V}(G,S)$. Since $T_{k,V}(G,S)$ is finite, $\ker h$ is torsion, and any $C]\not\in\ker h$ clearly cannot be torsion since for any $n\in\mathbb{N}$, $n\cdot h_{C}$ outputs a nonzero value.
@@
under $h$ forms a diagonal basis of $CF(G,p)=CF(G)$. Explicitly, setting $C_{P}:=k[G/P]\to k$, we have
$h_{C_{P}}(H)=\begin{cases}1&H\leq P\\ 0&H\not\leq P\end{cases}$
and a Möbius inversion argument shows this is a $\mathbb{Z}$-basis for $CF_{b}(G)$. Hence $h$ is surjective, and in fact, the above set is a basis for $\mathcal{E}_{k}^{V(\mathcal{F}_{G})}(G)$. ∎
@@
It is clear that every $kG$-module that has an endosplit $p$-permutation resolution is endo-$p$-permutation, but, in general, not every endo-$p$-permutation module has an endosplit $p$-permutation resolution. Mazza proved in [Ma03] when $p$ is odd and $G$ is a $p$-group, every endopermutation $kG$-module (hence endo-$p$-permutation kGkG-module) has an endosplit $p$-permutation resolution. On the other hand, if $G$ is a $2$-group, not every endopermutation $kG$-module has an endosplit $p$-permutation resolution; for instance, the unique endotrivial $kQ_{8}$-module that does not descend to $\mathbb{F}_{2}$ has no endosplit $p$-permutation resolution.
@@
    $p$-permutation resolution $C$ as follows. Let $\mathcal{X}_{C}$ denote the set of all $p$-subgroups of $G$ for which $C(P)\not\simeq 0$. It is easy to see that $\mathcal{X}_{C}$ forms a downward closed sub-poset of $s_{p}(G)$ which is closed under $G$-conjugation. For each $P\in\mathcal{X}_{C}$, write $h_{C}(P)$ for the unique integer for which $H_{h_{C}(P)}(C(P))\neq 0$. This defines a superclass function on $\mathcal{X}_{C}$, $h_{C}\colon\mathcal{X}_{C}\to\mathbb{Z}$.
@@
    We have that $C_{M}$ is a $V(\mathcal{F}_{G})$-endotrivial chain complex; we say a complex homotopy equivalent to a shift of a complex in the above form is a chain tensor of relative syzygies. Given any chain tensor of relative syzygies $C$, there exists a unique corresponding module tensor of relative syzygies, obtained by taking the nonzero homology $H_{h_{C}(1)}(C)$. Conversely, it is clear from the definition that given any module $M$ tensor of relative syzygies, there exists some chain tensor of relative syzygies $C_{M}$ resolving $M$, however $C_{M}$ need not be unique (even up to contractibles).
@@
We are about to characterize all $kG$-modules which have endosplit $p$-permutation resolutions, as well as all endosplit $p$-permutation resolutions themselves. This may be seen as the closest one may get to a classification theorem in the spirit of Dade, as there is no group structure that parameterizes all endo-$p$-permutation kGkG-modules (which have endosplit $p$-permutation resolutions) or all endosplit $p$-permutation resolutions, due to a lack of invertibility of some objects, in a sense.
Similar to how capped endopermutation modules can be thought of as the invertible endopermutation modules, and thus can be given a group structure, the $V(\mathcal{F}_{G})$-endotrivial complexes can be thought of as the invertible, or “capped,” endosplit $p$-permutation resolutions. The following theorem morally states that every endo-$p$-permutation module (resp. endosplit $p$-permutation resolution) can be built from invertible endo-$p$-permutation modules (resp. endosplit-$p$-permutation resolutions), similar to how all endopermutation modules for $p$-groups arise from capped endopermutation modules (see [D78a, Theorem 6.6]).